OpenStudy (anonymous):

2x+5y=10 solve for y

OpenStudy (anonymous):

2x + 5y = 10 5y = 10 - 2x 5y = 2(5-x) y = ?

OpenStudy (anonymous):

2x+5y=10 -2x -2x 5y=-2x+10 y=-2/5x+2?

OpenStudy (anonymous):

woah i just had this exact same problem earlier! :O but for me i had to find the intercepts for y and x. Answer: one is 2 the other is 5

OpenStudy (anonymous):

typo earlier: i meant one answer is -2 the other is 5. figure it out.